### v1

# Improving Code Reviewer Recommendation: Accuracy, Latency, Workload, and Bystanders

Abstract.
Code review ensures that a peer engineer manually examines the code before it is integrated and released into production. At Meta, we develop a wide range of software at scale, from social networking to software development infrastructure, such as calendar and meeting tools to continuous integration. We are constantly improving our code review system, and in this work we describe a series of experiments that were conducted across 10’s of thousands of engineers and 100’s of thousands of reviews.
We build upon the recommender that has been in production since 2018, RevRecV1. We found that reviewers were being assigned based on prior authorship of files. We reviewed the literature for successful features and experimented with them with RevRecV2 in production. The most important feature in our new model was the familiarity of the author and reviewer, we saw an overall improvement in accuracy of 14 percentage points.
Prior research has shown that reviewer workload is skewed. To balance workload, we divide the reviewer score from RevRecV2 by each candidate reviewers workload. We experimented with multiple types of workload to develop RevRecWL. We find that reranking candidate reviewers by workload often leads to a reviewers with lower workload being selected by authors.
The bystander effect can occur when a team of reviewers is assigned the review. We mitigate the bystander effect by randomly assigning one of the recommended reviewers. Having an individual who is responsible for the review, reduces the time take for reviews by -11%.
1. Introduction
Code review is critical activity in the software development lifecycle that ensures that at least one other experienced developer examines the code for defects and other issues. Code review has evolved from the inspection of massive completed artifacts (Fagan, 1976) to the continuous code review of the difference between the changed files in the form of pull-requests, patchsets, and diffs (Rigby and Bird, 2013; Bacchelli and Bird, 2013).
Expertise has been repeatedly shown to be a key factor in good code reviews (Porter et al., 1998; Rigby et al., 2014). Reviewer recommenders have been developed to suggest candidate reviewers who are most likely to perform high quality reviews, e.g., (Balachandran, 2013; Thongtanunam et al., 2015). Much of this work has been based on experience, e.g., this person has reviewed and authored the files in a change in the past. Other types of features including the relationships among developers have been used to identify candidate reviewers, e.g., (Jiang et al., 2017; Zhang et al., 2022). Most works have only tried to solve the problem of finding an expert reviewer, however, recent works have started to consider other factors such as knowledge spreading (Mirsaeedi and Rigby, 2020) and reviewer workload (Asthana et al., 2019; Strand et al., 2020; Al-Zubaidi et al., 2020)
At Meta, we are constantly evolving our code review system. In this work, we describe a series of improvements that we made to our code reviewer recommenders. Our starting point was developed in 2018 and we refer to it as RevRecV1. To address latency and accuracy issues with RevRecV1, we developed RevRecV2 that is based on a literature review of successfully used features. Building on RevRecV2, we use measures of reviewer workload to spread review work and develop RevRecWL. Finally, we combat the bystander effect when a group is assigned to do a review with BystanderRecRnd. One of our major contributions is to experiment with our recommenders in production and to conduct A vs B double blind trials on each recommender across hundreds of thousands of reviews and 10’s of thousands of engineers. We provide an overview of the motivation for each recommender and briefly summarize the results.
1. Accuracy and latency improvements with RevRecV2. On the code review team’s feedback group, some engineers complained that they were being incorrectly recommended for reviews. When we investigated, we found that the current recommender, RevRecV1, was purely based on hg blame lines and would suggest reviewers who had moved onto other projects because they had edited the files in the past. We reviewed the literature and developed a recommender that took the most promising features into account.
Result summary: In an A/B test, we saw a top 3 accuracy of 73% a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at the 90th percentile. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. In Summer 2022, RevRecV2 was rolled out to 100% of engineers.
2. Reviewer workload balancing with RevRecWL. The number of reviews that each engineer does is known to be skewed (Rigby et al., 2014; Hajari, 2022; Al-Zubaidi et al., 2020). At Microsoft (Asthana et al., 2019) and Ericsson (Strand et al., 2020), they implemented a recommender that weighted the expertise based recommender score by the workload of the candidate reviewer. We replicate these works and experiment with novel workload measures, including the number of scheduled meetings to understand if workload balancing is effective.
Result summary: When a reasonable candidate with lower workload was available, we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in review cycle time or latency. In Winter of 2023, RevRecWL was rolled out to 100% of engineers.
3. Reducing the bystander effect with BystanderRecRnd. The bystander effect occurs when nobody is explicitly assigned responsibility for a situation and no individual takes action or they are delayed in taking action (Fischer et al., 2011). Beyond noting that it may be an issue in code review (Meneely et al., 2014), we did not find any research in the software engineering literature. At Meta teams and groups of developers can be assigned a review without any explicit individual. To reduce the bystander effect, we use our recommenders to explicitly assign a reviewer.
Result summary: At Meta, we randomly assigned one of the top3 experts to review the diff and found a drop of -11.6% in time-in-review with no regressions in our guardrail metrics. In Spring of 2023, we rolled BystanderRecRnd out to 100% of engineers.
This works is structured contributions:
  1. (1)
    In Section 2, we describe Meta and its review process.
  2. (2)
    In Section 3, we review the literature to identify the features that have been successfully used to recommend reviewers.
  3. (3)
    In Section 5, we update the the recommender at Meta to include these features and experiment with a new model in production.
  4. (4)
    In Section 6, we weight the recommender by the workload of the candidate reviewers in an attempt to balance workload.
  5. (5)
    In Section 7, we reduce the impact of bystander effect when no individual is assigned a review by randomly assigning one of the recommended reviewers
  6. (6)
    We conduct A/B tests on each recommender including 100’s of thousands of reviews and 10’s of thousands of engineers across a diverse range of software from web infrastructure and social networking to internal tooling including continuous integration systems.
In the final sections of the paper, we discuss threats to validity in Section 8, position our work in the literature in Section 9, and conclude the paper in Section 10.
2. Background and Code Review Process
Figure 1. A code change, i.e. diff, under review at Meta. The recommended candidates for the review are clickable by the author and ordered left to right with the top ranked candidate on the left (see the red box in the figure). We have anonymized the diff and used mock names.
Meta is a large online services company that works across the spectrum in the communications, retail, entertainment industries. Meta has tens of thousands of employees with offices spread across the globe (North America, Europe, Middle East and Asia). Meta has its own dedicated infrastructure teams where the tools used are a mix of commercial and in-house developed systems. We study code reviews from the projects using Phabricator  [^1] Phabricator is an open source project available at https://www.phacility.com/phabricator/ which includes user facing social network products and virtual and augmented reality projects as well as software engineering infrastructure, such as calendar, tasks, and release engineering tooling.
2.1. Code review process
Code review at Meta follows a similar structure to the contemporary practices used at Google (Sadowski et al., 2018) and Microsoft (Bacchelli and Bird, 2013), and we briefly summarize the practice here. Figure 1 provides an example of a diff review with portions redacted and using mock names. Like most major code review tools, the code author provides a title and a detailed summary of the change. A test plan is provided that details which steps the author took to validate their changes, and provides reviewers a guide to test the change themselves. Line-by-line highlighting is used to show additions and removals, and inline comments are used to indicate any compilation or linting issues. The diff will also show the output of tests that were required to run based on the files included in the change.
Before publishing a code change, authors select one or more reviewers and/or reviewer groups. Reviewers provide feedback, ask questions and make suggestions, which should be addressed by the author. There may be multiple rounds of feedback and changes before a reviewer accepts the diff. Meta uses a single track/monorepo development model, so once the diff is accepted it will be integrated for further testing and ultimately “shipped“ into production.
For this paper, we focus on the reviewer recommendation experience. We see the recommended reviewers highlighted with a red box in Figure 1. We adapt the approaches behind the recommendations, but leave the interface consistent and unchanged.
2.2. Reviewer Recommenders at Meta
The reviewer recommender used in production at Meta before we began our work was developed in 2018. For this study it will be our starting comparison point and we will refer to it as RevRecV1. RevRecV1 considered the following attributes: the blame lines i.e. the person who has authored the most lines in a file is an expert, authorship frequency i.e. the person who has made the most changes to a file is an expert, and review frequency i.e. the number of times a reviewer has accepted the changes to a file. RevRecV1 is trained using a genetic algorithm, an ML approach popular in 2018. While the recommender has been effective and makes thousands of recommendations every day, the code review team identified a number of issues with it. First, many developers complain about incorrect recommendations on the feedback group. Second, hg-blame is an computationally expensive command that can delay the loading of the code review page for a diff. For these reasons, the code review team decided to test alternative strategies for recommending reviewers to improve accuracy and reduce latency.
3. Literature and Features
We reviewed the literature to identify the reviewer recommendation approaches and model features that were most successful in existing recommenders. Like Microsoft (Asthana et al., 2019; Zanjani et al., 2016) and Ericsson (Strand et al., 2020), we cannot use features in production that would add latency to our recommender. We also found that very few recommenders that are studied in the literature report latency and even fewer are actually deployed into production. We break the literature into the following categories (Al-Zubaidi et al., 2020): code ownership, reviewing experience, review participation rate, and familiarity with the author. Table 2 shows the features we selected for RevRecV2.
Code Ownership
The influence of code ownership on code quality has been extensively investigated in the literature(Bird et al., 2011; Rahman and Devanbu, 2011; Foucault et al., 2015; Thongtanunam et al., 2016) with recommenders to identify expert developers for a particular part of code or task (Mockus and Herbsleb, 2002). Researchers have used a wide range of granularity, from lines (Fritz et al., 2007; Girba et al., 2005; Rahman and Devanbu, 2011) to modules (Bird et al., 2011), to estimate ownership of developers. Studies on code review find that code owners are usually selected to review changes (Sadowski et al., 2018; Bacchelli and Bird, 2013; Greiler et al., 2016). In this work, we quantify ownership as the number of files under review that a candidate reviewer has modified in the past diffs.
Reviewing Experience
Thongtanunam et al. (Thongtanunam et al., 2016) recommend reviewers based on the files that an engineer has reviewed. Intuitively, engineers who have reviewed the files in the diff in the past, will be good candidate reviewers. Zanjani et al. (Zanjani et al., 2016) build upon this notion of reviewer expertise by adding the number of comments a candidate reviewer has made on the files under review. In this work, we count the number of comments as well as other reviewer actions such as accepting or rejecting the diff.
Familiarity with Patch Author
Moving beyond code ownership, reviewer recommenders have begun to consider relationships among the reviewer and the author (Bosu et al., 2016). The simplest metric is the number of times a reviewer and has commented on an authors diffs in the past. More sophisticated usages include social relationship graphs (Yu et al., 2016). With 10’s of thousands of developers, and thousands of diffs every day, social network graphs are not performant. In this work, we suggest a novel usage of how long a reviewer looks at an authors diffs: the time a reviewer spent looking at the diffs the author wrote in the code review tool. While Meta has sophisticated tracking of how long an engineer is looking at a webpage and uses this to capture how long a reviewer looks at a review, other projects can easily add this to the review tool. For example, the SmartBear code review tool has been tracking this information since 2006 (Cohen, 2006) to identify reviewers that looked at the change for too short a period of time given the size of the diff. The web abounds with code to monitor time spent on a webpage. [^2] A JavaScript library that tracks time spent on a page https://stackoverflow.com/questions/4667068/how-to-measure-a-time-spent-on-a-page
Review Participation Rate
On open source software projects, the number of reviews that a developer has been assigned and participated in is a strong predictor of whether they will perform a review (Ruangwan et al., 2019; Rigby et al., 2014). We measure this participation rate in two ways, the number of diffs that a reviewer was assigned and the number of diffs that a reviewer resigned from. These two measures capture how likely a reviewer is to participate in reviewing.
Workload
Recent works have incorporated the reviewer workload into reviewer recommendation. Early works quantified the relationship between time waiting for a review and review workload. Baysal et al. (Baysal et al., 2013) studied WebKit to understand the impact of metrics on review time. They found that when reviewers had longer queues of waiting reviews, reviews tended to take longer. They also found that more active reviewers had quicker response times than less active reviewers. Studying four open source projects, Ruangwan et al. (Ruangwan et al., 2019) found prior participation and prior experience were the best predictor of whether a reviewer will respond to a review request. The familiarity of the author an reviewer had a minor impact on participation.
Microsoft weighted their code reviewer recommender, cHRev, by the number of open reviews a candidate, WhoDo, had and found between a 14% and 21% improvement in the time taken for reviews depending on the project (Asthana et al., 2019). In contrast, Strand  et al. (Strand et al., 2020) evaluate a workload aware recommender at Ericsson. In a backtest they found that a recommender aware of the file change history and the level of recent workload activity was able to make recommendations with an MRR of .37 or an average rank of 2.7. Despite these reasonable recommendations, after manually evaluating the performance of the recommender on 47 changes they determined that it did not reduce lead time or workload for code review and they decided to not rollout the recommender.
4. Recommender Design and Evaluation Metrics
Table 1. We report accuracy, and we also have goal and guardrail metrics which we discuss for each experiment.
Outcome Metric Description
TopN Accuracy
How often does one of the recommended reviewers review the diff? We display 5 reviewers and report Top1 and Top3 accuracy.
TimeInReview
The time when a diff is waiting for review or is being actively reviewed. It excludes the time when the author is reworking the code. We want to make sure that our recommender does not slow down reviewers.
TimeSpent
The amount of time reviewers spend examining each diff. This is time spent on the review page for a diff. We want to make sure that reviewers are not rushing through reviews and rubberstamping diffs.
Latency
The amount of time from when the diff is published until the recommendations are displayed. We monitor the 90th and 99th percentile as our goal is to reduce the time it takes for an author to see candidate reviewers
Clicks
The number of times that an author selects one of the recommended reviewers.
Workload
We measure the amount of time a candidate reviewer has spent looking at diffs in the last week.
Table 2. The Features for RevRecV2 sorted by the feature importance in the model deployed in Section 5. We only include features that can be calculated with low latency, we use the feature categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020).
Category Description Feature Importance
Familiarity with Author
The time the candidate spent in the diff review tool looking at the author’s diffs in the past 90 days
33.89%
Review Participation Rate
Number of diffs the candidate was assigned as reviewer. At Meta reviewers are assigned to many diff either directly or as part of a team or review group.
31.28%
Reviewing Experience
Number of diffs the candidate rejected
11.56%
Reviewing Experience
Number of diffs the candidate commented or acted on as a reviewer
10.10%
Familiarity with Author
The time the candidate spent in the diff review tool looking at the author’s diffs when they were explicitly assigned as a reviewer in the past 90 days
6.69%
Code Ownership
Number of diffs the candidate authored in the past which touched any of the files in the diff.
4.91%
Reviewing Experience
Number of diffs the candidate commented on
1.57%
Review Participation Rate
Number of diffs the candidate resigned from. When a reviewer does not feel qualified to review a diff they can resign from that review.
0.00%
Having reviewed the literature on code review recommendation, we created a recommender that would suggest the reviewer that was most likely to perform a review, i.e. make a comment or do an action such as accepting a diff. Table 2 describes each of the features that we included across the categories of metrics found in the literature: code ownership, reviewing experience, reviewer participation, and familiarity with the author. We first develop a recommender considers all the prior features but excludes workload because the results of factoring workload are mixed, i.e. it worked at Microsoft but did not work at Ericsson. Instead, we weight the top ranked reviewers by their workload to understand if the approach that worked at Microsoft works at Meta.
4.1. Design for RevRecV2
Prior works have mostly ranked reviewers based on a set of features and rules, e.g., (Asthana et al., 2019; Mirsaeedi and Rigby, 2020). Since reviewer recommendation is a ranking problem, we select learn-to-rank approach (Burges, 2010), LambdaMART. A pairwise comparison of each reviewer is made and this ordering is then used to determine the final order. In our case we use MART (Multiple Additive Regression Trees) as function to determine the ordering of reviewers. At Meta there is infrastructural support for LambdaMART models at scale using measures collecting data from internal databases on code review and coding.
To train RevRecV2 we used the features shown in Table 2. We trained the model on over 4.3 million instances of these features for reviewers. We evaluated the model on 1.1 million recommended candidates. The feature importance can be seen in Table 2. We discuss A/B test experiment outcomes in Section 5.1 reporting both accuracy and latency results.
4.2. Design for RevRecWL
We design RevRecWL, which adds Workload (WL) to RevRecV2. Our goal is to identify the top expert developer with the lowest current workload. To this end, we adapt the approach taken at Microsoft (Asthana et al., 2019) that balances the workload of reviewers by weighting the score by the current workload of candidate reviewers for a diff review, $r$:
(1) $\text{\text{Load}(r)}=e^{\theta.Workload}$
$\theta$ is a parameter between 0 and 1 to control the amount of load balancing.
At Microsoft (Asthana et al., 2019), they found that initial feedback from developers indicated that reviewers from other projects complained that they were being incorrectly recommended. We made three compromises to convince the engineering team to run the experiment: (1) we set $0.1$ which reduces the influence of workload favoring expertise. (2) we only re-rank the top 5 suggestions from RevRecV2. (3) instead of assigning the top ranked reviewer, we allow the author to select one of the top5 reviewers. These have two important guaranties, first, people with low workload, but little experience will not be inadvertently added to the list eliminating the issue that Microsoft experienced. Second, the Top5 accuracy of RevRecWL = RevRecV2. Since the top 5 developers are displayed Phabricator, this means in the most extreme case, the top developer from RevRecV2 will be displayed as the 5th developer if they have high workload, but the same reviewers will be displayed only the order will change.
For a diff review, $r$, RevRecWL final scoring function is
(2) $\text{{RevRecWL}}(r)=\frac{\text{{RevRecV2}}(r)}{\text{Workload}(r)}$
Instead of choosing a single measure of workload, we experiment with three types of workload weights, two are novel: The number of upcoming meetings in the next 7 days, the reviewer activity in the last 7 days (Asthana et al., 2019), and time spent reviewing in the last 7 days. We calculate each over a 7 day period because developer patterns on a particular day tend to vary, but the entire work week tends to be more predictable (Chen et al., 2022). In Section 6, we present our historical analysis to determine the best workload measure and our A/B test to determine if we can use RevRecWL in production.
4.3. Design for RecBystander
The bystander effect is a well known psychological problem that when a group of people are present, it is less likely that any individual will take action (Fischer et al., 2011). At Meta, developers work in a monorepo, so there are times when the author does not know which engineer to assign the review to. In other cases, the author might not mind and anyone in on a particular team can do the review. We decided to use the recommenders we developed in the previous two sections to automatically assign a reviewer.
We experimented with two recommenders in production. BystanderRecRnd randomly assigns one of the top 3 reviewers from RevRecV2 to the review. BystanderRecWL builds on RevRecWL and takes the top reviewer with the a low workload and assigns them to the review. In both cases, we expected to see a decrease in the time-in-review. The results are presented in Section 7.
4.4. Goal and Guardrail Outcomes Metrics
The vast majority of reviewer recommendation work does not release the recommender in production and only evaluate the accuracy on on historical data. In contrast, the culture at Meta uses a sophisticated A/B test of every new feature and change. The team defines goal and guardrail metrics in advance of the experiment and then a double-blind randomized trail is used to test the existing feature, control, with the new feature, test. Goal metrics are the metrics we hypothesize will improve with the new feature. In contrast, guardrail metrics are the outcome metrics we want do not want to negatively impact. In Table 1, we introduce the the metrics that we use as outcomes in this paper: TopN accuracy, time-in-review, TimeSpent, Latency, Clicks on recommendations, and Workload. Depending on the experiment these outcomes will either be goal metrics or guardrails. For example, in the first experiment with RevRecV2 our goal is to improve TopN accuracy and reduce latency. Our guard metrics are to ensure that reviews do not rubberstamp reviews and rush them, i.e. TimeSpent does not decrease. We discuss the our goals and guardrails metrics in the experimentation sections below.
5. RevRecV2 in Production
In Section 3, we examined the code reviewer recommender literature. We use features from each category of measure we identified in the literature: code ownership, reviewer experience, reviewer participation rate, and familiarity with the author. While prior work have used how often a candidate reviewer in the past, we use a novel measure of how long an reviewer has viewed the author’s changes in the last 90 days TimeSpent.
We conducted an A/B test experiment to understand if the new model is effective in production. Our experimental unit was per diff review. Diffs in the control group, A, continued to receive reviewer recommendations from RevRecV1, the existing model in production. Diffs in the test group, B, received reviewer recommendations from our new model. Our experiment ran between in the late Spring of 2022, on 82k diff which were evenly and randomly split between the two conditions.
Our goal metrics, the metrics we wanted to improve, where the top 1, top 2 and top 3 recommendation accuracy. TopN accuracy in this context is the fraction of times where at least one person belonging to the top n predictions has either commented or acted, e.g., accepted, the diff. We also wanted to improve the latency on the slowest diffs and tracked the latency at the 90th and 99th percentiles.
Our guardrail metrics, the metrics we did not want to negatively influence, were the time diffs were in review (cycle time) and the TimeSpent that reviewers looked at each diff. We did not want to increase cycle time and we did not want recommended engineers to rubberstamp diffs and spend less time looking at them.
5.1. Experimental results for RevRecV2
In Table 3, we see that the new model is much more accurate than RevRecV1. Looking at top-1 accuracy, top-2 accuracy and top-3 accuracy we see 9.14, 12.53 and 14.19 percentage point improvements, respectively, over RevRecV1. In Table 2, we show the feature importance. The most important features that contribute to this accuracy improvement relate to the familiarity of the author with the review (TimeSpent on the author’s prior diffs, 33.89%) and the reviewer participation rate (number of diffs assigned 31.28%). In contrast, code and review ownership are still useful but much less important. The number of files in the current diff that the candidate reviewer has authored in the past accounts for only 4.91%. The number of diffs that a reviewer has acted on and rejected have an importance of 10.10% and 11.56%. Our findings clearly show that being a more active reviewer and being often reviewing the authors diffs are much stronger predictors of who will actually perform a review than code ownership.
Table 3. Improvement in Accuracy for RevRecV2
Top-1 Top-2 Top-3
RevRecV1 43.40 54.36 59.05
RevRecV2 52.54 66.89 73.24
Improvement 9.14 points 12.53 points 14.19 points
When comparing latency of the new model to RevRecV1, we see major wins in Table 4. The 90th percentile latency decreased from 4.43s to 0.3s (a 14x improvement), and p99 latency decreased from 12.37s to 1.26s (a 9x improvement). This reduction in latency largely because our new model uses measures that are stored in a database and does not require blame information to be calculated on the fly.
Table 4. Improvement in Latency for RevRecV2. We measure how long it takes for the recommender to make a recommendation in production, i.e. the latency. We are interested in the longest latency, so we report the 90th and 95th percentile, i.e. p90 and p95.
Latency at p90 Latency at p99
RevRecV1 (control) 4.43 s 12.37 s
RevRecV2 (test) 300 ms 1.26 s
Improvement 14x improvement 9x improvement
We also see that diffs in the test condition have more usage/clicks of the suggested reviewer feature, with a 27.3% increase in the number of diffs where a suggested employee reviewer was added using the recommender reviewer component on the diff review page, see Table 5. As well as a 9% increase in in the number of recommendations made. This is partially a result of higher accuracy predictions leading to more clicks as well as lower latency allowing authors to see the recommendations sooner.
Based on these results, RevRecV2 has been rolled out to 100% of engineers at Meta since early Summer 2022.
Table 5. We report the number of diffs that the models were able to create recommendations for and how often the author felt that they were useful, i.e. clicked the recommended reviewer.
Diff with recs Clicked recommendation
RevRecV1 (control) 38,247 6,420
RevRecV2 (test) 41,692 8,179
Improvement 9.00% 27.4%
In the A/B test, we saw a top 3 accuracy of 73% a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at p90. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. RevRecV2 has been rolled out to 100% of engineers at Meta.
6. Balancing Reviewer Workload
Table 6. Backtest comparison of metrics and results from workload balancing in production. We report the percentage point decrease in accuracy and the percentage decrease in workload relative to results from RevRecV2.
Condition Workload Metric Top1 Accuracy Top1 Workload Top3 Accuracy Top3 Workload
Backtest Upcoming Meetings -21.66 -1.46 -12.66 1.55
Backtest Reviewer Activity -20.61 -57.11 -13.40 -30.03
Backtest Time Spent -21.37 -72.79 -20.89 -41.14
Production Time Spent -4.80 -18.06 -5.34 -2.84
Engineers have different skillsets which will will always skew reviewer workload (Rigby and Storey, 2011), it is still interesting to understand if a recommender that is aware of workload might distribute the load of reviewing and potentially drive down cycle time by suggesting the least busy expert. For example, Microsoft weighted their code reviewer recommender by a candidate’s workload and reduced the time it takes for a review (Asthana et al., 2019). A similar experiment at Ericsson (Strand et al., 2020) balanced reviewer workload, but was unsuccessful and was rolled back. As a result, we build upon the Microsoft work and we take the score from RevRecV2 that is now in production and weight it by the workload of a candidate as described in Section 4.2.
We experiment with three types of workload weights the number of upcoming meetings in the next 7 days, the reviewer activity in the last 7 days (Asthana et al., 2019), and time spent reviewing in the last 7 days. We calculate each over a 7 day period because developer patterns on a particular day tend to vary, but the entire work week tends to be more predictable (Chen et al., 2022). We conduct a historical backtest to determine which measure works as well as a workload weight. We then roll out a workload in production and determine if it is effective.
Our goal metric is to reduce the median amount of time spent reviewing of the reviewer who made an action on a diff. We expect a decrease in accuracy, however, because we only re-rank the top 5 reviewers, the same reviewers will be shown they will just be ordered by their workload, so top5 precision will be unchanged.
Our other guardrail metrics are the same is in the previous section. We do not not want to increase review cycle time, we do more quickly reviewed diffs, so the time spent per diff should remain unchanged, we do not want to see fewer overall clicks on our recommendations, and we do not want latency to increase.
6.1. Historical Backtest
Most of the code review recommender literature uses historical data to simulate how well the recommender would have done at recommending the person who actually does the review (Thongtanunam et al., 2015; Mirsaeedi and Rigby, 2020). This accuracy is often measured with TopN accuracy. We also measure the median workload of the top recommended candidate. The top portion of Table 6, shows the results from 1.2 million historical diff reviews that were performed early 2023. We see that the Top1 accuracy decreases dramatically using workload balancing by around -20% compared to RevRecV2. The workload differences are less consistent, with upcoming meetings performing poorly, and review activity and time spent performing much better with respective decreases of -57% and -73%. The Top3 accuracy shows similar patterns with time spent having the largest drop in workload.
6.2. RevRecWL in Production
Since we are only re-ordering the reviewers suggested by RevRecV2 and are still displaying the Top5 recommended reviewers, we decided to use the most aggressive weight metric, TimeSpent7d, because it had the largest drop in workload. The unit in the experiment was the author of a diff, so each other would either have RevRecV2 or RevRecWL. We had 14k diff authors in each condition for a total of 28k authors.
We were surprised how different the results from the backtest were from the actual people who performed the review, see the last row in Table 6. When the author selected the top recommendation, the accuracy decreased much less than in the back test -5% vs -21% decrease in accuracy and the workload only decreased -18% vs -73% in the production vs backtest. A similar discrepancy can be seen with Top3, recall that Top5 will be unchanged.
Our guardrail metric, number of clicks explains the discrepancy, we saw a decrease in selection of the Top1 candidate by 30% of all diffs to 20%. Overall, however, the difference in the number of clicks did not change in a statistically significant manner. What this means is that when someone with low workload was an inappropriate reviewer, the author simply ignored them and moved lower in the list of recommended reviewers. This results also reinforces how important it is to first model expert reviewers with RevRecV2 and then simply re-rank them.
The remaining guardrail metrics latency, time spent per diff, and review cycle time where not changed in a statistically significant manner. We were unable to replicate the Microsoft (Asthana et al., 2019) improvement in review interval.
When a reasonable candidate with lower workload was available we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in review cycle time or latency. As a result, we decided to roll RevRecWL out to 100% at Meta.
7. The Bystander Effect
When a task is assigned to a group of people, the likelihood of any individual acting is lower than when an individual is explicitly assigned to this task (Fischer et al., 2011). At Meta, diff authors can assign teams and groups of developers rather than selecting individual reviewers. In some cases, an author might not mind which developer on a team reviews a diff. Furthermore, Meta uses a monorepo where developers often make changes that crosscut the codebase and are owned by another development team. The author might not know who the best reviewer is and will instead select a team. Selecting a reviewer group creates the risk of each member thinking, “someone else will probably review this.” We decided to leverage the wins from our reviewer recommendation investments last year to assign a single individual reviewer to diffs where only reviewer groups were assigned in an attempt to reduce the Bystander Effect.
We developed and tested two strategies to explicitly assign a reviewer on diffs where the author did not choose an individual reviewer. (1) BystanderRecWL, automatically assigns the top candidate reviewer from RevRecWL to the diff, i.e. the top expert with the current lowest workload. (2) BystanderRecRnd, the simpler of the two, built upon RevRecV2 and simply chose one of the top three reviewers at random. Our goal is that reviewers will feel a sense of ownership for diffs where they were assigned as a reviewer, and review it more quickly, thus reducing the time-in-review metric. We do not want to see regressions in how long a developer views a diff. We do not track latency, clicks, or workload as they are unlikely to be impacted by this experiment.
7.1. Results for Bystander Effect
We had expected BystanderRecWL to outperform the simpler
BystanderRecRnd because we figured that a low workload person would respond more quickly. We were surprised to find that BystanderRecWL had no statistically significant impact on any of the metrics including our goal metric of reducing time-in-review.
In contrast, BystanderRecRnd, was extremely successful and reduced time-in-review by -11.6%. There was no statistically significant change in the guardrail metric of the amount of time reviewers spent on each diff. We rolled BystanderRecRnd out to 100% of diffs at Meta.
When a new feature or change is rolled out at Meta, we post an announcement in our feedback groups. In this case, we had one developer voice concern that explicitly assigning an individual to a diff will take time away from other diffs. To investigate this concern, we looked at all diffs during the time of study and saw no statistically significant change in the time in review per diff for diffs that already had an individual assigned to the diff.
The bystander effect occurs in code review when only a team is assigned to a review a diff instead of an individual. At Meta we randomly assigned one of the top3 experts to review the diff and found a drop of -11.6% in time-in-review with no regressions in our guardrail metrics. We rolled BystanderRecRnd out to 100% of engineers.
8. Threats to Validity
8.1. Generalizability
Drawing general conclusions from empirical studies in software engineering is difficult because any process depends on a potentially large number of relevant context variables. The analyses in the present paper were performed at Meta, and it is possible that results might not hold true elsewhere. We cannot release our data, even in an anonymized format, because it would violate legally binding employee privacy agreements. However, our study does cover a very wide swath of software engineering. The software systems covers millions of lines of code and 10’s of thousands of developers who are both collocated and working at multiple locations across the world. We also cover a wide range of domains from user facing social network products and virtual and augmented reality projects to software engineering infrastructure, such as calendar, task, and release engineering tooling.
We have discussed the many reviewer recommenders available in the literature above. Instead of reimplementing these recommenders, we have used the feature categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020), and included features from each categories that would not increase latency in our model. It is possible that smaller organizations or projects can use code review social network features etc, unfortunately they do not work at our scale.
8.2. Construct Validity
We have also reviewed the literature to ensure that we covered the normal features/constructs include in reviewer recommenders. We have used those that can be extracted from Meta’s history and those that can be calculated in a timely manner.
The implicit assumption when using historical data is that the person who performs an action on a diff is a good reviewer. Recommenders have been trained on this dataset (Kovalenko et al., 2018). However, as we saw in, for example, our RevRecWL evaluation, when recommenders that have alternative purposes, i.e. spreading workload, the results can be drastically different in a historical backtest compared to an actual release in production. To ensure that we cover a wide range of constructs, we use multiple goal and guardrail measures in each experiment and do not just rely on results from a backtest.
8.3. Internal Validity
Each of the recommenders presented in this paper have been tested in a production experiment based on goal and guardrail metrics. However, each experiment was run at a different time, with RevRecV2 experimentation in the Spring of 2022, RevRecWL in winter of 2023, and BystanderRecRnd in Spring of 2023. We cannot report the exact time-frames, but report the number of diffs or diff authors so that the size of sample can be evaluated by the reader. As part of engineering at Meta we track metrics across time and did not see any changes in accuracy, time-in-review, and TimeSpent outside the experiments that we ran.
9. Literature and Discussion
The first reviewer recommendation system we are aware of was introduced by Balachandran et al. (Balachandran, 2013). They used authorship of the changed lines in a code review (using git blame) to identify who had worked on that code before and suggested a ranked list of this set as potential reviewers. Similarly, Thongtanunam et al. (Thongtanunam et al., 2015) proposed RevFinder, a reviewer recommender based on file locations. RevFinder is able to recommend reviewers for new files based on reviews of files that have similar paths in the filesystem. The approach was evaluated on over 40,000 code reviews across three OSS projects, and recalls a correct reviewer in the top 10 recommendations 79% of the time on average. Like RevRecV1 these approaches would suggest reviewers who had moved onto other projects because they had edited the files in the past.
There has been a lot of interest in consuming features beyond file history when building reviewer recommendation models. Rahman et al. (Rahman et al., 2016) propose CORRECT an approach to recommend reviewers based on their history across all of GitHub as well as their experience certain specialized technologies associated with a pull request. Jiang et al. (Jiang et al., 2017) examine the impact of various attributes of a pull request on a reviewer recommender, including file similarity, PR text similarity, social relations, and “activeness” and time difference. They find that adding measures of activeness to prior models increases performance considerably. Zhang et al. (Zhang et al., 2022) built a Socio-Technical graph to capture the social interactions among authors and reviewers. Further, they use Graph Convolution Networks (GCNs) to encode the social interaction information into high-dimensional space, and formulate reviewer recommendation as a link prediction problem. We describe the features that we were able to include in our model in Table 2 building upon the categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020).
Latency and Performance. Some of these recommenders use social network features and other features that required substantial processing. As noted by Microsoft (Asthana et al., 2019; Zanjani et al., 2016) and Ericsson (Strand et al., 2020), these cannot be used in production systems due to processing delay. However, they did not report results on the importance of latency. Interestingly, when we reduced the latency from RevRecV1 to RevRecV2 we saw that lower latency resulted in authors seeing recommendations earlier before they had decided on who would do the review and many more clicks on recommendations. Future works should provide information on prediction latency.
9.1. Reviewer Workload and Recommenders
Kovalenko et al. (Kovalenko et al., 2018) found that expertise based recommenders tend to suggest reviewers that are always obvious to authors 52% and 63% of the time at Microsoft and JetBrains respectively. They are usually obvious to authors 37% and 38% of the time. As a result, recommendations often clicked on because it is easier to click than type a reviewer’s name. Another interesting finding was that authors often found that outdated recommendations were made based on people who use to work on the files. Authors wanted more information about the current knowledge level of the candidate reviewers as well as how busy, i.e. what the workload of the reviewers are.
Microsoft weighted their code reviewer recommender, cHRev, by the number of open reviews a candidate, WhoDo, had and found between a 14% and 21% improvement in the time taken for reviews depending on the project (Asthana et al., 2019). We were unable to replicate this result at Meta and found no statistically significant change in review cycle time. We did find a decrease in workload of selected reviewers. Unfortunately, Microsoft did not report the change in reviewer workload, so we are unable to make comparisons.
Al-Zubaidi et al. (Al-Zubaidi et al., 2020) treat reviewer recommendation as a multi-objective search problem where they maximize for the chance of participating in review and minimze the shannon entropy, i.e. skewness, of the people selected. Their recommender considers review workload, but has low precision ranging from 16% to 20% depending on the open source project. They do not use a measure of workload as an outcome, so it is unclear if their approach succeeds in spreading workload. They do not report on latency or performance, and it is unclear if this search-based approach could be used at large scale companies.
Strand  et al. (Strand et al., 2020) evaluate a workload aware recommender at Ericsson. In a backtest they found that a recommender aware of the file change history and the level of recent workload activity was able to make recommendations with an MRR of .37 or an average rank of 2.7. Despite these reasonable recommendations, after manually evaluating the performance of the recommender on 47 changes they determined that it did not reduce lead time or workload for code review and they decided to not rollout the recommender.
Murphy-Hill et al. (Murphy-Hill et al., 2023), examine workload at Google and discover an systemic imbalance in gender and code review selection. Authors manually select men more often than women as reviewers. Since recommenders are trained on prior data, the recommenders contain this bias. In our work, we do not explicitly factor gender into our recommender, however, we re-rank reviewers based on workload, so if women are assigned fewer reviewers, they will automatically be assigned more reviews as workload is balanced. We also use random in our bystander experiment, which Murphy-Hill et al. finds balances the gender.
10. Concluding Remarks
We make the following contributions related to accuracy and latency, workload, and the bystander effect:
  1. (1)
    RevRecV2: In the A/B test, we saw a top 3 accuracy of 73% a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at p90. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. RevRecV2 has been rolled out to 100% of engineers at Meta.
  2. (2)
    RevRecWL: When a reasonable candidate with lower workload was available we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in review cycle time or latency. As a result, we decided to only rerank the top three reviewers, and rolled RevRecWL out to 100% at Meta.
  3. (3)
    BystanderRecRnd: The bystander effect occurs in code review when only a team is assigned to a review a diff instead of an individual. At Meta we randomly assigned one of the top3 experts to review the diff and found a drop of -11.6% in time-in-review with no regressions in our guardrail metrics. We rolled BystanderRecRnd out to 100% of engineers.
We hope that our works will inspire other engineers, researchers, and companies to look beyond experience based recommendation to consider author-reviewer familiarity, reviewer workload, and the bystander effect of only recommending teams of reviewers.